OpenStudy (anonymous):

Find the area. a. 18.98 cm2 c. 9.9 cm2 b. 9.49 cm2 d. 19.8 cm2

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

@milk333

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

9.49

OpenStudy (anonymous):

7.3 * 2.6 =18.98 then you have divide by 2 and get 9.49

OpenStudy (anonymous):

b

OpenStudy (anonymous):

fan or medal me please

OpenStudy (anonymous):

A= 1/2 * b*h (Base times height) A= 1/2* 2.6*7.3 (Convert 1/2 into a decimal, 0.5) A= 0.5*2.6*7.3 A= 9.49